PA9-3 Part 1-b to 3 1-b. Prepare the journal entries for each of the above transactions.

January 2 Purchased a bulldozer for $250,000 by paying $20,000 cash and signing a $230,000 note due in five years.

January 3 Replaced the steel tracks on the bulldozer at a cost of $20,000, purchased on account. The new steel tracks increase the bulldozer's operating efficiency.

January 30 Wrote a check for the amount owed on account for the work completed on January 3. February 1 Repaired the leather seat on the bulldozer and wrote a check for the full $800 cost. March 1 Paid $3,600 cash for the rights to use computer software for a two-year period.


PA9-3 Part 1-b to 3


1-b. Prepare the journal entries for each of the above transactions.

2. For the tangible and intangible assets acquired in the preceding transactions, determine the amount of depreciation and amortization that Precision Construction should report for the quarter ended March 31. The equipment is depreciated using the double-declining-balance method with a useful life of five years and $40,000 residual value.

3. Prepare a journal entry to record the depreciation and amortization calculated in requirement 2 .
Answer #1

1.

Date Account Titles Debit Credit
Jan-02 Bulldozer $       250,000
       Cash $          20,000
       Note Payable $       230,000
(Purchased bulldozer)
Jan-03 Bulldozer $          20,000
        Accounts Payable $          20,000
(Replaced tracks on bulldozer)
Jan-30 Accounts Payable $          20,000
      Cash $          20,000
(Paid cash)
Feb-01 Repair and Maintenance Expense $                800
      Cash $                800
(Repaired seat of bulldozer)
Mar-01 Computer Software $            3,600
      Cash $            3,600
(Purchase computer software)

2.
Double Declining Depreciation = Beginning Book Value x 2 times straight line rate
Straight line rate = 1/5 x 100 = 20%

Depreciation (3 months) = $270000 x 40% x 3/12 = $27000
Amortization (1 month) = $3600 / 2 x 1 / 12 = $150

3.

Date Account Titles Debit Credit
Mar-31 Depreciation Expense $          27,000
       Accumulated Depreciation $          27,000
(Depreciation recorded)
Mar-31 Amortization Expense $                150
      Computer Software $                150
(Software amortized)